OpenStudy (anonymous):

if f(x)=2(squared)x& g(x)=x^2-1 which statement is true?

OpenStudy (anonymous):

A) (fog)(x)=2(x^2-1) B) (fog)(x)=2x(x^2-1) C) (gof)(x)=4x^2-1 D) (gof)(x)=4x-1

OpenStudy (anonymous):

@Fall12-13

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@Alfonso92

OpenStudy (anonymous):

\[F(X)=\sqrt{X}, G(X)=X ^{2}-1\]

OpenStudy (anonymous):

Soooo, A?